The molar solubility of aluminum hydroxide in a 0.179 M aluminum acetate solution is




M.

1
Expert's answer
2022-04-11T04:23:05-0400

Al(CH3COO)3 = Al3+ + 3CH3COO-

[Al3+] = 0.179 mol

Ksp = [Al3+] × [OH-]3 = (0.179)(3S)3 = 4.833S3

S3 = Ksp / 4.833 = 1.3 × 10-33 / 4.833 = 2.7 × 10-32

S = 6.5 × 10-12 mol/L